# Automating Steering for Safe Multimodal Large Language Models

### v1

@@ Abstract
Recent progress in Multimodal Large Language Models (MLLMs) has unlocked powerful cross-modal reasoning abilities, but also raised new safety concerns, particularly when faced with adversarial multimodal inputs. To improve the safety of MLLMs during inference, we introduce a modular and adaptive inference-time intervention technology, AutoSteer, without requiring any fine-tuning of the underlying model. AutoSteer incorporates three core components: (1) a novel Safety Awareness Score (SAS) that automatically identifies the most safety-relevant distinctions among the model’s internal layers; (2) an adaptive safety prober trained to estimate the likelihood of toxic outputs from intermediate representations; and (3) a lightweight Refusal Head that selectively intervenes to modulate generation when safety risks are detected. Experiments on LLaVA-OV and Chameleon across diverse safety-critical benchmarks demonstrate that AutoSteer significantly reduces the Attack Success Rate (ASR) for textual, visual, and cross-modal threats, while maintaining general abilities. These findings position AutoSteer as a practical, interpretable, and effective framework for safer deployment of multimodal AI systems.
@@
Lyucheng Wu1, Mengru Wang1,2, Ziwen Xu1,2, Tri Cao3, Nay Oo3, Bryan Hooi3, Shumin Deng3thanks: Corresponding Author. 1Zhejiang University   2Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph 3National University of Singapore, NUS-NCS Joint Lab, Singapore {lyuchengwu,mengruwg,ziwen.xu}@zju.edu.cn{dcsbhk,shumin}@nus.edu.sg
@@ 1 Introduction
Recent advances in Multimodal Large Language Models (MLLMs) Wu et al. (2023); Yin et al. (2024) have enabled models to understand, reason, and generate across modalities. With unified Transformer architectures and efficient alignment techniques, MLLMs excel in tasks like visual question answering, multimodal dialogue, and grounded generation (Team, 2024; Li et al., 2025), driving their adoption in real-world domains such as education, healthcare, and content creation.
@@
Despite these advancements, the growing power of MLLMs also raises pressing safety concerns. Prior work has shown that MLLMs are susceptible to generating harmful, offensive, or unethical content, particularly when triggered by adversarial inputs from either the textual or visual modality (Han et al., 2025; Chen et al., 2024). Compared to unimodal language models, MLLMs present unique safety challenges due to their rich multimodal input space and the potential for harmful outputs to emerge from subtle cross-modal interactions. Furthermore, harmfulness is often nuanced and context-dependent, resisting binary classification and complicating the design of static safety filters. Overly aggressive safety interventions risk suppressing benign outputs or degrading overall model utility, highlighting the need for more precise and adaptive safety alignment techniques.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) process and reason over data from multiple modalities—such as text, images, audio, and video—within a unified architecture, typically based on Transformers (Wu et al., 2023). Formally, the input is a sequence of modality-tagged tokens:
$\mathcal{X}=\{x_{1}^{(m_{1})},x_{2}^{(m_{2})},\dots,x_{n}^{(m_{n})}\},\quad m_{i}\in\mathcal{M}$
@@
Existing MLLMs fall into two main categories: modality-encoder-based and early-fusion. The former uses separate encoders $E_{m}$ per modality, aligns features via fusion modules (e.g., cross-attention), and decodes through a language model. For example, BLIP-2 Li et al. (2023) combines a frozen vision encoder, a lightweight Query Transformer, and an LLM:
$\mathcal{Y}=f_{\theta}(\textsf{QFormer}(E_{\text{img}}(\text{Img}));\text{Text})$
@@
$\mathcal{Y}=f_{\theta}(\mathcal{X})$
@@ 2.2 Steering Techniques
Model Steering, also referred to as controllable text generation (CTG) (Liang et al., 2024), aims to guide large language models (LLMs) to produce outputs that exhibit desired attributes such as sentiment, style, or safety. Among various techniques, Latent Space Manipulation (LSM) has emerged as a flexible and model-agnostic approach. Rather than modifying model parameters or architecture, LSM operates directly on the model’s internal representations, typically word embeddings or hidden states, at inference time. A notable example is LM-Steer (Han et al., 2024), which introduces a simple but effective mechanism to steer generation by linearly transforming the output embeddings of a frozen language model (LM). Let $\mathbf{c}\in\mathbb{R}^{d}$ denote the context vector at a given decoding step, in standard decoding, the probability of generating the token $v$ is calculated by:
@@ 3.1 Overview
The framework consists of three components: (1) layer selection to identify an internal layer $L$ with salient safety signals; (2) a safety prober $\mathcal{P}$ trained to detect unsafe content from $L$’s representations; and (3) a conditional refusal head that modulates generation based on $\mathcal{P}$’s confidence.
@@ 3.2 Safety-Aware Layer Selection
which captures the representational shift indicative of safety versus toxicity distinctions, while controlling for confounding linguistic factors such as syntax and named entities.
@@
where $\cos(\cdot,\cdot)$ denotes cosine similarity. A higher SAS score indicates stronger alignment among contrastive vectors, suggesting that the corresponding layer encodes safety-relevant distinctions in a more consistent and structured manner. The layer with the highest SAS score is selected as the target for subsequent safety probing, which will be $L$.
@@ 3.3 Safety Prober
We train a safety prober $\mathcal{P}$ (utilizing 3,000 toxic examples from the VLSafe dataset and 3,000 constructed safe examples, as detailed in Section 4.1) to distinguish between safe and toxic inputs using the activation vectors extracted from the selected layer ${L}$. Specifically, for a given input $x$, we denote its activation at layer ${L}$ as $h_{{L}}(x)$. The prober $\mathcal{P}$ maps this activation to a scalar score via:
@@ 3.4 Automatic Adaptive Steering
The output score $s=\mathcal{P}(h(x))$ from the prober is then used to control the model’s generation behavior. To implement this, we define a steering signal $\alpha(s)$ that determines the degree to which the model should be steered away from generating unsafe responses. This steering signal is mapped through a thresholding function:
@@
where $\tau$ is a threshold and can be adjusted according to need. Normally, $\tau$ is set to 0.5. This function ensures adaptive application of steering.
We integrate this control mechanism into the generation process using the LM-Steer framework. Specifically, let $\mathbf{e}_{v}$ denote the output embedding of a vocabulary token, $\mathbb{I}$ be an identity matrix, $\epsilon\in\mathbb{R}$ be a scalar that controls the polarity and intensity of the steering effect and $\mathbf{W}$ be the learned Refusal Head matrix (trained on 3,000 toxic samples from the VLSafe dataset, referring to Section 4.1 for details). The steered embedding $\mathbf{e}^{\prime}_{v}$ is computed as:
@@
where $\mathbf{H}_{t}$ represents the hidden state at decoding step $t$, which serves as the input to the output embedding layer for computing the logits. $\mathbf{E}$ denotes the original output embedding matrix consisting of token embeddings $\mathbf{e}_{v}$, and $\mathbf{E}^{\prime}$ represents the modified output embedding matrix composed of LM-Steer-transformed token embeddings $\mathbf{e}^{\prime}_{v}$, which replaces $\mathbf{E}$ during decoding.
The final token probabilities are derived from $\tilde{z}_{t}$, thereby modulating the generation output according to the safety signal. This mechanism ensures that safety constraints are imposed only when necessary, thereby maintaining the model’s fluency and overall generative capabilities.
@@
Table 1: Attack Success Rate (ASR, %) and Accuracy (Acc, %) of MLLMs under a steer intensity of $\epsilon=0.1$. Lower ASR ($\downarrow$) indicates better safety , while higher Acc ($\uparrow$) reflects better general utility. MMMU results on Chameleon are omitted due to performance being near random guess levels. The difference from results on original (Orig.) base model is caculated in the bracket. As shown, AutoSteer performs well: (1) On Llava-OV, it achieves similarly low ASR as the Steer baseline, indicating high safety, while fully preserving the utility of Orig. model. (2) On Chameleon, AutoSteer also demonstrates strong performance, except on the ToViLaG+ test set containing only toxic images, where the globally low Safety Awareness Score of Chameleon hinders effective prober training.
@@ 4.3 Overall Performance
On Chameleon, AutoSteer also delivers substantial ASR reductions across all benchmarks: On VLSafe, ASR is reduced from 67.8% to 15.4%, matching the Steer baseline; On ToViLaG+, AutoSteer outperforms the original model by large margins across all subsets. While Steer marginally outperforms AutoSteer on the image-only subset, this exception likely results from Chameleon’s lower safety awareness at the representation level, limiting the prober’s ability to detect image-driven toxicity. This is further supported by Figure 4 and Appendix E.1, where the SAS and prober accuracy on Chameleon lags behind Llava-OV. Importantly, even under this challenging condition, AutoSteer achieves a substantial ASR drop of 8.3 percentage points, offering robust safety gains with minimal overhead and without requiring aggressive interventions that could impair utility.
@@
On Llava-OV, AutoSteer matches or slightly exceeds both the original model and Steer: On RealWorldQA, AutoSteer attains 61.8% accuracy, on par with the original model (61.8%) and outperforming Steer (60.8%); On MMMU, AutoSteer achieves 48.4%, again slightly higher than Steer (47.8%). On Chameleon, despite the base model’s relatively limited capacity, AutoSteer still matches the original performance (6.0%) and surpasses Steer (5.4%) on RealWorldQA. These results validate that AutoSteer’s conditional design preserves utility by default. Instead of uniformly steering all inputs, it selectively triggers detoxification only when toxicity is detected. This avoids unnecessary interference with benign queries, preventing the degradation commonly observed in standard steer-based approaches.
@@ 4.4.1 Evaluation of SAS
To locate safety-relevant layers, we train a separate prober on each layer and evaluate its ability to distinguish toxic from non-toxic inputs. Training uses a dedicated validation set, and testing is conducted on safe (RealWorldQA) and toxic (VLSafe and ToViLaG+) benchmarks. We show Accuracy trends across training epochs on LLaVA-OV in Figure 3, and SAS scores for LLaVA-OV and Chameleon in Figure 4. Full results are in Appendix E.1.
For LLaVA-OV, most layers achieve high probing accuracy (>98%) on both training and testing sets. Layer 20 consistently outperforms other layers, aligning with its high SAS score shown in Figure 4. Interestingly, early layers (e.g., 4 and 8) achieve good performance on text-only and text-image-both toxicity subsets but fail on image-only ones, contradicting their SAS rankings, suggesting shallow correlation rather than deep safety awareness. Besides, only mid-to-late layers (16 and 20) generalize well to image-only toxicity subsets, aligning well with their SAS scores.
For Chameleon, early layers (4, 8, 12) exhibit weak probing ability (shown in Figure 10 at Appendix E.1) and low SAS scores (shown in Figure 4). In contrast, mid-to-late layers (16, 20) achieve good performance (>90% accuracy) on text-only and text-image-both toxicity subsets, again corresponding with higher SAS. Notably, no layer is effective on the image-only toxicity subset, reflecting Chameleon’s limited capacity to encode visual safety concepts. Together, these findings validate SAS as a reliable metric for selecting prober layers, especially in multimodal safety scenarios.
@@ 4.4.2 Mechanism Behind SAS
4.4.3 Can the Prober’s Toxicity Score Truely Reflect Toxicity?
@@ 4.4.4 Effect of Steering Intensity $\epsilon$ on ASR
Importantly, this plateau effect introduces a practical trade-off. While higher $\epsilon$ can detoxify more effectively, it may also increase the risk of unintended consequences, such as semantic drift or degraded general performance (see Section 4.4.5). Thus, selecting an appropriate $\epsilon$ is crucial: too low may fail to mitigate toxic responses; too high may unnecessarily alter benign outputs.
@@ 5 Related Work
Steering techniques aim to control model behavior, such as sentiment, style, or safety, while preserving coherence. Training-phase methods include control-specific architectures Keskar et al. (2019); Zhang et al. (2020); Hua and Wang (2020), lightweight tuning Zeldes et al. (2020); Zhou et al. (2023), and RL-based optimization Upadhyay et al. (2022); Ouyang et al. (2022); Dai et al. (2024). Inference-phase methods include prompt-based steering Shin et al. (2020); Li and Liang (2021), latent-space steering Liu et al. (2024b); Chan et al. (2021), and decoding-time control Dathathri et al. (2020); Krause et al. (2021). These techniques adjust model outputs in real-time without requiring retraining.
@@
Inference-time safety defense is a promising approach to enhance MLLM safety without retraining. CoCA Gao et al. (2024) adjusts token logits based on safety prompts, while ECSO Gou et al. (2024) transforms unsafe images into safer text captions. InferAligner Wang et al. (2024) uses cross-model guidance to improve safety during inference, and Immune Ghosal et al. (2024) formalizes defense mechanisms as a decoding problem. These methods effectively reduce toxic outputs while maintaining model utility, making them practical solutions for real-time safety enhancement in MLLMs.

### v3

@@ Abstract
Recent progress in Multimodal Large Language Models (MLLMs) has unlocked powerful cross-modal reasoning abilities, but also raised new safety concerns, particularly when faced with adversarial multimodal inputs. To improve the safety of MLLMs during inference, we introduce a modular and adaptive inference-time intervention technology, AutoSteer, without requiring any fine-tuning of the underlying model. AutoSteer incorporates three core components: (1) a novel Safety Awareness Score (SAS) that automatically identifies the most safety-relevant distinctions among the model’s internal layers; (2) an adaptive safety prober trained to estimate the likelihood of toxic outputs from intermediate representations; and (3) a lightweight Refusal Head that selectively intervenes to modulate generation when safety risks are detected. Experiments on LLaVA-OV and Chameleon across diverse safety-critical benchmarks demonstrate that AutoSteer significantly reduces the Attack Success Rate (ASR) for textual, visual, and cross-modal threats, while maintaining general abilities. These findings position AutoSteer as a practical, interpretable, and effective framework for safer deployment of multimodal AI systems [^1] The code: https://github.com/zjunlp/AutoSteer..
@@
Lyucheng Wu1,3, Mengru Wang1,2, Ziwen Xu1,2, Tri Cao3, Nay Oo3, Bryan Hooi3, Shumin Deng3thanks: Corresponding Author. 1Zhejiang University   2Zhejiang University - Ant Group Joint Lab of Knowledge Graph 3National University of Singapore, NUS-NCS Joint Lab, Singapore lyuchengwu@u.nus.edu231sm@zju.edu.cn{dcsbhk,shumin}@nus.edu.sg
@@ 1 Introduction
Recent advances in Multimodal Large Language Models (MLLMs) Wu et al. (2023); Yin et al. (2024) have enabled models to understand, reason, and generate across modalities. With unified Transformer architectures and efficient alignment techniques, MLLMs excel in tasks like visual question answering, multimodal dialogue, and grounded generation (Team, 2024; Li et al., 2025), driving their adoption in real-world domains such as education and healthcare Song et al. (2025).
@@
Despite these advancements, the growing power of MLLMs also raises pressing safety concerns. Prior work has shown that MLLMs are susceptible to generating harmful, offensive, or unethical content, particularly when triggered by adversarial inputs from either the textual or visual modality (Han et al., 2025; Chen et al., 2024). Compared to unimodal language models, MLLMs present unique safety challenges due to their rich multimodal input space and the potential for harmful outputs to emerge from subtle cross-modal interactions. Furthermore, harmfulness is often nuanced and context-dependent, complicating binary classification and the design of static safety filters. Overly aggressive safety interventions risk suppressing benign outputs or degrading overall model utility, highlighting the need for more precise and adaptive safety alignment techniques.
@@ 2.1 Multimodal Large Language Models
Multimodal Large Language Models (MLLMs) process and reason over data from multiple modalities ($\mathcal{M}$), such as text, images, audio, and video, within a unified architecture, typically based on Transformers (Wu et al., 2023). Formally, the input is a sequence of modality-tagged tokens:
$\mathcal{X}=\{x_{1}^{(m_{1})},x_{2}^{(m_{2})},\dots,x_{n}^{(m_{n})}\},\quad m_{i}\in\mathcal{M}$ (1)
@@
Existing MLLMs fall into two main categories: modality-encoder-based and early-fusion. The modality-encoder-based models use separate encoders $E_{m}$ per modality, aligns features via fusion modules (e.g., cross-attention), and decodes through a language model. For example, BLIP-2 Li et al. (2023) combines a frozen vision encoder, a lightweight Query Transformer, and an LLM:
$\mathcal{Y}=f_{\theta}(\textsf{QFormer}(E_{\text{img}}(\text{Img}));E_{\text{txt}}(\text{Text}))$ (2)
@@
$\mathcal{Y}=f_{\theta}(\mathcal{X})$ (3)
@@ 2.2 Steering Techniques
Model Steering, also referred to as controllable text generation (CTG) (Liang et al., 2024), aims to guide large language models (LLMs) to produce outputs that exhibit desired attributes such as sentiment, style, or safety. Among various techniques, Latent Space Manipulation (LSM) has emerged as a flexible and model-agnostic approach. Rather than modifying model parameters or architecture, LSM operates directly on the model’s internal representations, typically word embeddings or hidden states, at inference time. A notable example is LM-Steer (Han et al., 2024), which introduces a simple but effective mechanism to steer generation, by linearly transforming the output embeddings of a frozen language model (LM). Let $\mathbf{c}\in\mathbb{R}^{d}$ denote the context vector at a given decoding step, in standard decoding, the probability of generating the token $v$ is calculated by:
@@ 3.1 Overview
As shown in Figure 2, AutoSteer framework consists of three components: (1) layer selection learned to identify an internal layer $L$ with salient safety signals; (2) a safety prober $\mathcal{P}$ trained to detect unsafe content from $L$’s representations; and (3) a conditional refusal head that modulates generation based on $\mathcal{P}$’s confidence.
@@ 3.2 Safety-Aware Layer Selection
which captures the representational shift indication of safety versus toxicity distinctions, while controlling for confounding linguistic factors such as syntax and named entities.
@@
where $\cos(\cdot,\cdot)$ denotes cosine similarity. A higher SAS score indicates stronger alignment among contrastive vectors, suggesting that the corresponding layer encodes safety-relevant distinctions in a more consistent and structured manner. The layer with the highest SAS score is selected as the target for subsequent safety probing, denoted as $L$.
@@ 3.3 Safety Prober
We train a safety prober $\mathcal{P}$ (utilizing 3,000 toxic examples from the VLSafe dataset and 3,000 constructed safe examples, as elaborated in §4.1) to distinguish safe and toxic inputs using the activation vectors extracted from the selected layer L{L}. Specifically, for a given input $x$, we denote its activation at layer L{L} as $h_{{L}}(x)$. The prober $\mathcal{P}$ maps this activation to a scalar score via:
@@ 3.4 Automatic Adaptive Steering
The output score $s=\mathcal{P}(h(x))$ from the prober is then used to control the model’s generation behavior. To achieve this, we define a steering signal $\alpha(s)$ that determines the degree to which the model should be steered away from generating unsafe responses. This steering signal is determined by a thresholding function:
@@
where $\tau$ is a tunable threshold, typically set to 0.5. This function ensures that steering is applied adaptively, activating only when the output score exceeds the specified threshold.
We integrate this control mechanism into the generation process using LM-Steer Han et al. (2024) framework. Specifically, given the output embedding $\mathbf{e}_{v}$ of a vocabulary token, an identity matrix $\mathbb{I}$, a scalar $\epsilon\in\mathbb{R}$ that controls the polarity and intensity of the steering effect, and the learned Refusal Head matrix $\mathbf{W}$ (trained on 3,000 toxic samples from the VLSafe dataset, referring to §4.1 for details), the steered embedding $\mathbf{e}^{\prime}_{v}$ is computed as:
@@
where $\mathbf{H}_{t}$ represents the hidden state at decoding step $t$, which serves as the input to the output embedding layer for computing the logits. $\mathbf{E}$ denotes the original output embedding matrix consisting of token embeddings $\mathbf{e}_{v}$, and $\mathbf{E}^{\prime}$ represents the modified output embedding matrix composed of token embeddings $\mathbf{e}^{\prime}_{v}$ transformed by LM-Steer, which replaces $\mathbf{E}$ during decoding.
The final token probabilities are derived from $\tilde{z}_{t}$, thereby modulating the generation output according to the safety signal. This mechanism ensures that safety constraints are imposed only when necessary, thereby maintaining the model’s fluency and general capabilities.
@@
Table 1: Attack Success Rate (ASR, %) and Accuracy (Acc, %) of MLLMs under a steer intensity of $\epsilon=0.1$. Lower ASR ($\downarrow$) indicates better safety , while higher Acc ($\uparrow$) reflects better general utility. MMMU results on Chameleon are omitted due to performance being near random guess levels. The difference from results on original (Orig.) base model is calcualted in the bracket. As shown, AutoSteer performs well: (1) On Llava-OV, it achieves similarly low ASR as the Steer baseline, indicating high safety, while fully preserving the utility of Orig. model. (2) On Chameleon, AutoSteer also demonstrates strong performance, except on the ToViLaG+ test set containing only toxic images, where the globally low Safety Awareness Score of Chameleon hinders effective prober training.
@@ 4.3 Overall Performance
On Chameleon, AutoSteer also delivers substantial ASR reductions across all benchmarks: On VLSafe, ASR is reduced from 67.8% to 15.4%, matching the Steer baseline; On ToViLaG+, AutoSteer outperforms the original model by large margins across all subsets. While Steer marginally outperforms AutoSteer on the image-only subset, this exception possibly results from Chameleon’s lower safety awareness at the representation level, limiting the prober’s ability to detect image-driven toxicity. This is further supported by Figure 4 and Appendix E.1, where the SAS and prober accuracy on Chameleon lags behind Llava-OV. Importantly, even under this challenging condition, AutoSteer achieves a substantial ASR drop of 8.3 percentage points, offering robust safety gains with minimal overhead and without requiring aggressive interventions that could impair utility.
@@
On Llava-OV, AutoSteer matches the original model and exceeds Steer: On RealWorldQA, AutoSteer attains 61.8% accuracy, on par with the original model (61.8%) and outperforming Steer (60.8%); On MMMU, AutoSteer achieves 48.4%, slightly higher than Steer (47.8%). On Chameleon, despite the base model’s relatively limited capacity, AutoSteer still matches the original model (6.0%) and surpasses Steer (5.4%) on RealWorldQA. These results validate that AutoSteer’s conditional design preserves utility by default. Instead of uniformly steering all inputs, it selectively triggers detoxification only when toxicity is detected. This avoids unnecessary interference with benign queries, preventing degradation commonly observed in standard steer-based approaches.
@@ 4.4.1 Evaluation of SAS
To locate safety-relevant layers, we train a separate prober on each layer and evaluate its ability to distinguish toxic from non-toxic inputs. Training uses a dedicated validation set, and testing is conducted on safe (RealWorldQA) and toxic (VLSafe and ToViLaG+) benchmarks. We show Accuracy trends across training epochs on LLaVA-OV in Figure 3, and SAS scores for LLaVA-OV and Chameleon in Figure 4. Full results are shown in Appendix E.1.
For LLaVA-OV, most layers achieve high probing accuracy (>98%) on both training and test sets. Layer 20 consistently outperforms other layers, aligning with its high SAS score shown in Figure 4. Interestingly, early layers (e.g., 4 and 8) achieve good performance on text-only and text-image-both toxicity subsets but fail on image-only toxicity ones, suggesting shallow correlation rather than deep safety awareness. Besides, only mid-to-late layers (16 and 20) generalize well to image-only toxicity subsets, aligning well with SAS scores.
For Chameleon, early layers (4, 8, 12) exhibit weak probing ability (as shown in Figure 10 at Appendix E.1) and low SAS scores (as shown in Figure 4). In contrast, mid-to-late layers (16, 20) achieve good performance (>90% accuracy) on text-only and text-image-both toxicity subsets, again corresponding with higher SAS scores. Notably, no layer is effective on the image-only toxicity subset, reflecting Chameleon’s limited capacity to understand visual safety concepts. Together, these findings demonstrate SAS as a reliable metric for selecting prober layers, especially in multimodal safety scenarios.
@@ 4.4.2 Mechanism Behind SAS
4.4.3 Can the Prober’s Toxicity Score Truly Reflect Toxicity?
@@ 4.4.4 Effect of Steering Intensity $\epsilon$ on ASR
Importantly, this plateau effect introduces a practical trade-off. While higher $\epsilon$ can detoxify more effectively, it may also increase the risk of unintended consequences, such as semantic drift or degraded general performance (referring to §4.4.5). Thus, it’s quite crucial to select an appropriate $\epsilon$: too low may fail to mitigate toxic responses; too high may unnecessarily alter benign outputs and damage model generation.
@@ 5 Related Work
Steering techniques aim to control model behavior, such as sentiment, style, or safety, while preserving coherence. This paradigm can be divided into two categories. Training-phase methods include control-specific architectures Keskar et al. (2019); Zhang et al. (2020); Hua and Wang (2020), lightweight tuning Zeldes et al. (2020); Zhou et al. (2023), and RL-based optimization Upadhyay et al. (2022); Ouyang et al. (2022); Dai et al. (2024). Inference-phase methods include prompt-based steering Shin et al. (2020); Li and Liang (2021), latent-space steering Liu et al. (2024b); Chan et al. (2021), and decoding-time control Dathathri et al. (2020); Krause et al. (2021). These techniques adjust model outputs in real-time without requiring retraining.
@@
Inference-time safety defense is a promising approach to enhance MLLM safety without retraining. CoCA Gao et al. (2024) adjusts token logits based on safety prompts, while ECSO Gou et al. (2024) transforms unsafe images into safer text captions. InferAligner Wang et al. (2024) uses cross-model guidance to improve safety during inference, and Immune Ghosal et al. (2024) formalizes defense mechanisms as a decoding problem. HiddenDetect Jiang et al. (2025) introduces a tuning-free approach to detect jailbreak attacks against MLLMs by monitoring refusal-related semantics in the internal activations. These methods effectively reduce toxic outputs while maintaining model utility, making them practical solutions for real-time safety enhancement in MLLMs.